# supplement

### v1

@@ 4 First steps: a Naive baseline
This allows us to compute the estimated miscoverage as a weighted sum for each quantile level $\tau$ via
$\hat{\alpha}(\tau)\;=\;\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\tau}(X_{i})\;\mathbb{I}\bigl{\{}\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr{\}}\;\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\tau}(X_{i})\bigr{\}}.$ (4)
@@
Armed with $\hat{\alpha}(\tau)$, we construct the calibrated LPB by selecting the largest $\tau$ such that $\hat{\alpha}(\tau)\leq\alpha$, i.e.,
$\hat{L}(x)=\hat{q}_{\hat{\tau}}(x),\qquad\hat{\tau}=\sup\bigl{\{}\tau\in\mathcal{T}:\sup_{\tau^{\prime}<\tau}\hat{\alpha}(\tau^{\prime})\leq\alpha\bigr{\}}.$
We summarize this Naive approach in Algorithm 1 in Appendix B.1 and analyze its validity in the next section.
4.1 PAC-based coverage validity and factors affecting its tightness
We build on the theory developed in [15] and adapt it to the setting where the true weights are known. In this setup, the Naive approach achieves a tighter coverage bound than the Adaptive-T procedure. The following proposition shows that the LPB $\hat{L}$ constructed by the Naive approach holds a PAC-type coverage guarantee.
Proposition 4.1 (Validity of the Naive calibration method).
@@ 4.1 PAC-based coverage validity and factors affecting its tightness
Fix a tolerance level $\delta\in\left(0,1\right)$ and a miscoverage level $\tau\in\left(0,1\right)$. Suppose that $\{(X_{i},T_{i})\}_{i\in{\mathcal{I}_{2}}}$ and $(X_{\textup{test}},T_{\textup{test}})$ are drawn i.i.d., $\hat{q}_{\tau}(x)$ is non-decreasing and continuous in $\tau$, and that there exists a constant ${\gamma}_{\tau}>0$ such that the weights from (3) satisfy ${w}_{\tau}(x)\leq{\gamma}_{\tau}$ for $P_{X}$-almost all $x$. Then, with probability at least $1-\delta$ over the draws of $\mathcal{D}$, the LPBs $\hat{L}(x)$ produced by Algorithm 1 satisfy
@@
The proof is given in Appendix D.6. Importantly, the above guarantee holds in finite samples, for any LLM $\mathcal{G}(x)$, any prompts distribution, and regardless of the accuracy of the quantile estimators $\hat{q}_{\tau}(x)$. However, this coverage bound depends on the supremum of all possible weights with non-zero probability, denoted by $\gamma_{\tau}$. In particular, as $\gamma_{\tau}$ increases, the bound becomes looser.
@@
In Appendix C, we further analyze the Naive method in greater depth by studying the properties of $\hat{\alpha}(\tau)$ defined in (4). Specifically, in Proposition C.2, we make an illustrative assumption that, for any fixed $\tau$, each time-to-unsafe-sampling $T_{i}$ is miscovered by $\hat{q}_{\tau}(X_{i})$ at the same constant rate once we condition on the calibration prompts $\{X_{j}\}_{j\in\mathcal{I}_{2}}$. Under this assumption, we prove that the conditional variance $\mathrm{Var}\!\bigl{[}\hat{\alpha}(\tau)\,\bigm{|}\,\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr{]}$ is linearly monotone with regard to the mean calibration weight $\overline{w}_{\tau}=\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in{\mathcal{I}_{2}}}\frac{1}{\pi_{i}}$. Notably, this result holds for all calibration methods proposed in this work.
The above discussion shows that any calibration scheme producing large weights will both weaken the PAC-type coverage bound (through a large supremum weight $\gamma_{\tau}$) and increase the variance of our miscoverage estimator (through a large mean weight $\bar{w}_{\tau}$). In the Naive calibration scheme, the event $\hat{q}_{\tau}(X_{i})\leq C_{i}$ can be rare, making $\bar{w}_{\tau}$ and $\gamma_{\tau}$ excessively large. Consequently, the resulting coverage rates would deviate from the nominal level—as reflected in our experiments in Section 6.
5 Advanced method: prompt-adaptive calibration
In this section, we build on the Naive method and present our main contribution, the Prompt-Adaptive Budget calibration algorithm. Our method operates in three modes, sequentially building upon each other: (1) the Basic mode makes the censoring times $C_{i}$ adaptive to the prompt $X_{i}$, enabling better utilization of the LLM sampling budget; (2) the Trimmed mode imposes an upper bound on $\gamma_{\tau}$ to tighten the coverage guarantee; and finally, (3) our flagship Optimized mode further refines the method by minimizing $\bar{w}_{\tau}$ to reduce the variance of the miscoverage estimator.
5.1 Basic mode: introducing prompt-adaptive sampling for improved budget utilization
@@
To better utilize the calibration budget in the computation of $\hat{\alpha}(\tau)$ in (4), we aim to design prompt-dependent censoring times $C_{i}$ that maximize $\mathbb{P}\bigl{(}\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr{)}$. However, before we can increase this probability, we first need to choose a quantile level $\tau$ to consider. For this purpose, we assume that the quantile estimators are monotone in $\tau$. This assumption can be enforced by sorting $\hat{q}_{\tau}(X_{i})$. Consequently, $\forall\tau_{1}\leq\tau_{2},\mathbb{I}\{\hat{q}_{\tau_{2}}(X_{i})\leq C_{i}\}\Rightarrow\mathbb{I}\{\hat{q}_{\tau_{1}}(X_{i})\leq C_{i}\}.$ Thus, setting $C_{i}$ to the 100%–quantile, i.e., $C_{i}=\hat{q}_{1}(X_{i})$, would maximize the effective calibration set for all quantile levels $\tau\in[0,1]$. However, in our case, $T_{i}$ follows a geometric distribution, and thus its 100%–quantile is infinite. As a result, this naive choice is infeasible under our finite budget constraint $\mathbb{E}\left[\sum_{i\in{\mathcal{I}_{2}}}C_{i}|\{X_{i}\}_{i\in{\mathcal{I}_{2}}}\right]\leq B$.
@@
Instead, we consider a more practical compromise: we select a quantile level $\tau_{\text{prior}}\ll 1$ that represents our prior belief $\hat{\tau}\in[0,\tau_{\text{prior}}]$. For example, if the target coverage level $1-\alpha$ is not too extreme (e.g., $90\%$) and the uncalibrated quantile estimator $\hat{q}_{\tau}$ is reasonably accurate, a prior level such as $\tau_{\text{prior}}=0.2$ would be a suitable choice. Importantly, constraining $\tau$ to lie within this prior range affects only the informativeness of the LPB, and not its validity; this is formally proven in Section 5.3.
After choosing the prior level $\tau_{\textup{prior}}$, we set each censoring threshold $C_{i}$ based on the estimated quantile $\hat{q}_{\tau_{\textup{prior}}}(X_{i})$. To ensure that the $i$-th sample contributes to $\hat{\alpha}(\tau)$ for every $\tau\in[0,\tau_{\textup{prior}}]$, $C_{i}$ must be at least $\hat{q}_{\tau_{\textup{prior}}}(X_{i})$. As such, we define the censoring times as
$C_{i}:=\text{Ber}(\pi_{i})\cdot\hat{q}_{\tau_{\text{prior}}}(X_{i}),$ (6)
where $\text{Ber}(\pi_{i})$ is a Bernoulli random variable with per-prompt evaluation probability
$\pi_{i}=\min({B}/({|{\mathcal{I}_{2}}|\cdot\hat{q}_{\tau_{\text{prior}}}(X_{i})}),1).$
This design of $C_{i}$ satisfies the budget constraint (1).
@@
Once all $C_{i}$-s are set, we employ the same calibration procedure introduced in Section 4, with the sole modification that the search space for $\tau$ is now $\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$. We remark that the per-prompt evaluation probability $\pi_{i}$ does not depend on $\tau$, and so the inverse-censoring weights become $w_{\tau}\bigl{(}X_{i}\bigr{)}=1/{\pi_{i}},$ for all $\tau\in\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$. Hence, we omit the dependence on $\tau$ in our notation and denote the weights as $w\bigl{(}X_{i}\bigr{)}$, the maximum weight by $\gamma$, and the average weight by $\bar{w}$. The full procedure is presented in Algorithm 2. In Section 5.3, we establish a lower bound on the coverage rate of this procedure. As with the Naive method, this bound tightens as $\gamma$ decreases. In the following section, we explicitly control the value of $\gamma$ by trimming the output of $\hat{q}_{\tau}(x)$.
5.2 Trimmed mode: tightening coverage by capping quantile estimates
@@
To limit the value of $\gamma$, we simply cap each estimated quantile at a fixed threshold $M\in\mathbb{N}$, defining
$\hat{f}_{\tau}(x):=\min(\hat{q}_{\tau}(x),M).$ (7)
Consequently, the censoring times now are given by
$C_{i}:=\text{Ber}(\pi_{i})\cdot\hat{f}_{\tau_{\text{prior}}}(X_{i}),$
where
$\pi_{i}=\min({B}/({|{\mathcal{I}_{2}}|\cdot\hat{f}_{\tau_{\text{prior}}}(X_{i})}),1).$
@@
The rest of the calibration process continues as outlined in Section 4, using the trimmed estimates $\hat{f}_{\tau}(X_{i})$ instead of $\hat{q}_{\tau}(X_{i})$; See Algorithm 2.111Note that now $\mathcal{T}=\bigl{\{}\sup_{\tau\in[0,1]}\{\hat{f}_{\tau}(x)\leq\tilde{T}_{i}\}:i\in{\mathcal{I}_{2}}\}\cup\{\sup_{\tau\in[0,1]}\{\hat{f}_{\tau}(x)\leq C_{i}\}:{i\in{\mathcal{I}_{2}}}\bigr{\}}\cup\{0\}$.
@@
Before proceeding to the Optimized method, we pause to discuss the implications of the threshold $M$ on the validity and power of the resulting LPBs. First, since $\hat{f}_{\tau_{\text{prior}}}(X_{i})\leq M$, all per-prompt evaluation probabilities satisfy $\pi_{i}\leq\min\left({B}/({|{\mathcal{I}_{2}}|\cdot M}),1\right)$, and so the maximum weight $w(x)$ is bounded by $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)$. As we will see in Theorem 5.2, this allows us to obtain a coverage guarantee that is closer to the desired $1-\alpha$ level. Second, the expression $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)$ offers guidance for choosing the sampling budget $B$: it highlights the trade-off between the tightness of the coverage guarantee and the power of the resulting LPBs. Specifically, one can increase $M$ while ensuring the value of $\gamma$ remains at a certain level (e.g., $\gamma=2$) by demanding a higher sampling budget $B$. Indeed, our experiments show that when the budget permits setting a sufficiently large $M$, trimming has a small effect on the power of the resulting LPBs. Third, from a practical perspective, trimming the quantile estimates limits the maximal number of LLM samplings per prompt, preventing any single prompt from demanding an infeasible computational or audit workload.
5.3 Optimized mode: distributing the sampling budget across prompts to reduce
variance

### v4

@@ 4 Calibration with a known censoring mechanism
Using these weights, we estimate the empirical miscoverage rate at quantile level $\tau$ as a weighted average over the calibration set:
$\hat{\alpha}(\tau)\;=\;\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\tau}(X_{i})\;\mathbb{I}\bigl\{\tilde{T}_{i}<\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr\}.$ (6)
Note that this estimator differs from the one used in (3): it does not require additional normalization since our proposed censoring mechanisms provide the true weights $w_{\tau}(X_{i})$, as explained in Section 5.
@@
Armed with $\hat{\alpha}(\tau)$, we obtain calibrated LPB by following (4) and selecting the largest $\tau$ such that $\hat{\alpha}(\tau)\leq\alpha$:
$\hat{L}(x)=\hat{q}_{\hat{\tau}}(x),\ \ \hat{\tau}=\sup\bigl\{\tau\in\mathcal{T}:\sup_{\tau^{\prime}<\tau}\hat{\alpha}(\tau^{\prime})\leq\alpha\bigr\},\ \ \$(7)
where $\mathcal{T}=\bigl\{\sup_{\tau\in[0,1]}\{\hat{q}_{\tau}(X_{i})\leq\tilde{T}_{i}\}:i\in{\mathcal{I}_{2}}\}\cup\{\sup_{\tau\in[0,1]}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}:{i\in{\mathcal{I}_{2}}}\bigr\}\cup\{0\}$.
The following informal theorem shows that the LPB $\hat{L}$ formulated in (7) holds a PAC-type coverage guarantee. In Appendix D.6 we state the formal version of this theorem and prove it by building on the theoretical framework developed by Gui et al. (2024).
Theorem 4.1 (General validity, informal).
@@
Fix a tolerance level $\delta\in\left(0,1\right)$ and a miscoverage level $\tau\in\left(0,1\right)$. Suppose that $\{(X_{i},T_{i})\}_{i=1}^{n}$ and $(X_{\textup{test}},T_{\textup{test}})$ are drawn i.i.d and that the censoring times satisfy the conditional independence assumption (Assumption 2.1). Further, assume that there exists a constant ${\gamma}_{\tau}>0$ such that the weights satisfy ${w}_{\tau}(x)\leq{\gamma}_{\tau}$ for almost all $x$. Then, with probability at least $1-\delta$ over the draws of $\mathcal{D}$, the LPB $\hat{L}(x)=\hat{q}_{\hat{\tau}}(x)$ from (7) satisfies
$\mathbb{P}\left[T_{\textup{test}}\geq\hat{L}(X_{\textup{test}})|\mathcal{D}\right]\geq 1-\alpha-\sup_{\tau\in[0,1]}\left\{\sqrt{\frac{2\gamma^{2}_{\tau}+5}{|{\mathcal{I}_{2}}|}\cdot\log\left(\frac{1}{\delta}\right)}\right\}.$
@@
Importantly, the above guarantee holds in finite samples, for any LLM $\mathcal{G}(x)$, any prompts distribution $P_{X}$, sampling-budget $B$, and regardless of the accuracy of the quantile estimators $\hat{q}_{\tau}(x)$. However, this coverage bound depends on the supremum of all possible weights with non-zero probability, denoted by $\gamma_{\tau}$. In particular, as $\gamma_{\tau}$ increases, the bound becomes looser.
@@
In Appendix C, we further analyze the calibration method in greater depth by studying the properties of $\hat{\alpha}(\tau)$ defined in (6). Specifically, in Proposition C.2, we make an illustrative assumption that, for any fixed $\tau$, each time-to-unsafe-sampling $T_{i}$ is miscovered by $\hat{q}_{\tau}(X_{i})$ at the same constant rate, conditional on the calibration prompts $\{X_{j}\}_{j\in\mathcal{I}_{2}}$. This assumption is satisfied by the oracle conditional quantile $q_{\tau}(\cdot)$ because, by definition, $\mathbb{P}\left[T_{i}<q_{\tau}(X_{i})\mid X_{i}\right]=\tau$ for every $i$, and hence also conditional on $\{X_{j}\}_{j\in\mathcal{I}_{2}}$. Under this assumption, we prove that the conditional variance $\mathrm{Var}\!\bigl[\hat{\alpha}(\tau)\,\bigm|\,\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr]$ is linearly monotone with regard to the mean calibration weight
$\overline{w}_{\tau}=\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in{\mathcal{I}_{2}}}w_{\tau}(X_{i}).$ (8)
The above discussion shows that any censoring design producing large weights will both weaken the PAC‐type coverage bound (through a large supremum weight $\gamma_{\tau}$) and increase the variance of the empirical miscoverage estimator (through a large mean weight $\bar{w}_{\tau}$). Indeed, these two properties guide our Optimized censoring design presented in the following section, where we both (i) cap $\gamma_{\tau}$ by a constant value; and (ii) minimize the mean weight $\bar{w}_{\tau}$ subject to fully utilizing the sampling-allocation budget $B$.
5 Proposed budget-allocation mechanism designs
Having formulated the general calibration scheme and the factors influencing its coverage guarantee, we now introduce a sequence of baseline sampling-budget allocation strategies—Naive, Basic, and Trimmed—that build up to our flagship Optimized approach.
5.1 Naive mode
As a warm-up, we introduce a simple, baseline budget allocation strategy, the Naive approach, which defines each censoring time $C_{i}$ as a random variable independent of the prompt $X_{i}$. Assuming $T_{i}|X_{i}$ is a Geometric random variable with infinite support, we design $C_{i}$ to be distributed geometrically as well: $C_{i}\sim\mathrm{Geom}(\min(|{\mathcal{I}_{2}}|/B,1)),\ \forall i\in{\mathcal{I}_{2}}$, which satisfies the budget constraint in (2) by design. This Naive approach is summarized in Algorithm 3 in Appendix B.
Under this Naive allocation, the event $\hat{q}_{\tau}(X_{i})\leq C_{i}$ can be rare, especially for large $\hat{q}_{\tau}(X_{i})$. This makes $\bar{w}_{\tau}$ and $\gamma_{\tau}$ excessively large. Consequently, the resulting coverage rates might deviate from the nominal level—as reflected in our experiments in Section 6. This discussion motivates the method presented below, which utilizes the budget more efficiently and designs the censoring time as a function of the prompt.
5.2 Basic mode
@@
To better utilize the calibration budget in the computation of $\hat{\alpha}(\tau)$ in (6), we aim to design prompt-dependent censoring times $C_{i}$ that maximize $\mathbb{P}\bigl(\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr)$. Ideally, we would like to set the censoring times for each $\tau$ to increase this probability. However, this is infeasible as we construct the dataset only once. As a result, we must commit to a single budget allocation by choosing a specific quantile level $\tau$ to consider. For this purpose, we assume that the quantile estimators are monotone in $\tau$. This assumption can be enforced by sorting $\hat{q}_{\tau}(X_{i})$. Consequently, $\forall\tau_{1}\leq\tau_{2},\ \mathbb{I}\{\hat{q}_{\tau_{2}}(X_{i})\leq C_{i}\}\Rightarrow\mathbb{I}\{\hat{q}_{\tau_{1}}(X_{i})\leq C_{i}\}.$ Thus, if we were to naively set $C_{i}=\hat{q}_{1}(X_{i})$ as the 100%–quantile, we would indeed maximize $\mathbb{P}\bigl(\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr)$ for all $\tau\in[0,1]$. However, $\hat{q}_{1}(X_{i})$ can be excessively large, e.g., this 100%-quantile is infinite if $T_{i}\mid X_{i}$ follows a geometric distribution. As a result, this naive choice is infeasible under our finite budget constraint $\mathbb{E}\left[\sum_{i\in{\mathcal{I}_{2}}}C_{i}|\{X_{i}\}_{i\in{\mathcal{I}_{2}}}\right]\leq B$.
@@
Instead, we consider a more practical compromise: we select a quantile level $\tau_{\text{prior}}\ll 1$ that represents our prior belief $\hat{\tau}\in[0,\tau_{\text{prior}}]$. For example, if the target coverage level $1-\alpha$ is not too extreme (e.g., $90\%$) and the uncalibrated quantile estimator $\hat{q}_{\tau}$ is reasonably accurate, a prior level such as $\tau_{\text{prior}}=20\%$ would be a suitable choice. Importantly, constraining $\tau$ to lie within this prior range affects only the informativeness of the LPB, and not its validity; see Appendix D.6.
After choosing the prior level $\tau_{\textup{prior}}$, we turn to set a prompt-adaptive censoring time $C_{i}$ using the estimated quantile $\hat{q}_{\tau_{\textup{prior}}}(X_{i})$. Following (6), the $i$-th sample contributes to $\hat{\alpha}(\tau)$ if $C_{i}\leq\hat{q}_{\tau_{\textup{prior}}}(X_{i})$ for all $\tau\in[0,\tau_{\textup{prior}}]$. Combining this observation with the monotonicity of $\hat{q}_{\tau}$, a sensible choice for $C_{i}$ that avoids wasted samplings would either be (i) $C_{i}=\hat{q}_{\tau_{\text{prior}}}(X_{i})$, or (ii) $C_{i}=0$. Other choices for $C_{i}$ that are strictly greater or smaller than $\hat{q}_{\tau_{\text{prior}}}(X_{i})$ would result in unnecessary generations that do not contribute to the miscoverage estimator $\hat{\alpha}(\tau)$. Motivated by this observation, we define $C_{i}:=\text{Ber}(\pi_{i})\cdot\hat{q}_{\tau_{\text{prior}}}(X_{i})$, where $\text{Ber}(\pi_{i})$ is a Bernoulli random variable with success probability $\pi_{i}=\min({B}/({|{\mathcal{I}_{2}}|\cdot\hat{q}_{\tau_{\text{prior}}}(X_{i})}),1).$ This design of $C_{i}$ satisfies the budget constraint (2).
@@
Once all $C_{i}$-s are set, we can use our calibration procedure from Section 4, with the sole modification that the search space for $\tau$ is now $\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$. We remark that the per‐prompt evaluation probability $\pi_{i}$ does not depend on $\tau$, and so the inverse‐censoring weights become $w_{\tau}\bigl(X_{i}\bigr)=1/{\pi_{i}},$ for all $\tau\in\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$. Hence, we omit the dependence on $\tau$ in our notation and denote the weights as $w\bigl(X_{i}\bigr)$, the maximum weight by $\gamma$, and the average weight by $\bar{w}$.
For reference, Algorithm 3 summarizes the calibration procedure using the above Basic budget allocation scheme. As with the Naive approach, the coverage bound for this method (Theorem 4.1) tightens as $\gamma$ and $\bar{w}$ decrease. In the following section, we explicitly control the value of $\gamma$ by trimming the output of $\hat{q}_{\tau}(x)$.
5.3 Trimmed mode
@@
To limit the value of $\gamma$, we simply cap each estimated quantile at a fixed threshold $M\in\mathbb{N}$, defining $\hat{f}_{\tau}(x):=\min(\hat{q}_{\tau}(x),M)$. Consequently, the censoring times are now given by $C_{i}:=\text{Ber}(\pi_{i})\cdot\hat{f}_{\tau_{\text{prior}}}(X_{i}),$ where
$\pi_{i}=\min({B}/({|{\mathcal{I}_{2}}|\cdot\hat{f}_{\tau_{\text{prior}}}(X_{i})}),1).$ (9)
@@
The rest of the calibration process continues as outlined in Section 4, using the trimmed estimates $\hat{f}_{\tau}(X_{i})$ instead of $\hat{q}_{\tau}(X_{i})$; See Algorithm 3.222Note that now $\mathcal{T}=\bigl\{\sup_{\tau\in[0,1]}\{\hat{f}_{\tau}(X_{i})\leq\tilde{T}_{i}\}:i\in{\mathcal{I}_{2}}\}\cup\{\sup_{\tau\in[0,1]}\{\hat{f}_{\tau}(X_{i})\leq C_{i}\}:{i\in{\mathcal{I}_{2}}}\bigr\}\cup\{0\}$ and the search space for $\tau$ is $\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$.
@@
Before proceeding to the Optimized method, we pause to discuss the implications of the threshold $M$ on the validity and statistical efficiency of the resulting LPBs. First, since $\hat{f}_{\tau_{\text{prior}}}(X_{i})\leq M$, then $\pi_{i}$ from (9) satisfies $\pi_{i}\leq\min\left({B}/({|{\mathcal{I}_{2}}|\cdot M}),1\right)$, and so the maximum weight $w(x)$ is bounded by $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)$. As observed in Theorem 4.1, this allows us to obtain a coverage guarantee that is closer to the desired $1-\alpha$ level. Second, the expression $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)$ offers guidance for choosing the sampling budget $B$: it highlights the trade-off between (i) the tightness of the coverage guarantee and (ii) the statistical efficiency of the resulting LPB. Specifically, one can increase $M$ while ensuring the value of $\gamma$ remains at a certain level (e.g., $\gamma=2$) by demanding a higher sampling budget $B$. Indeed, our experiments show that when the budget permits setting a sufficiently large $M$, trimming has a small effect on the power of the resulting LPBs. Third, from a practical perspective, trimming the quantile estimates limits the maximal number of LLM samplings per prompt, preventing any single prompt from demanding an infeasible computational or audit workload.
5.4 Optimized mode: our flagship method